OpenStudy (anonymous):

differentiate using limit sqrt (x+9)

OpenStudy (anonymous):

do you mean using the limit definition of a derivative?

OpenStudy (anonymous):

\[\lim_{h\to 0}\frac{\sqrt{x+h+9}-\sqrt{x+9}}{h}\] multiply top and bottom by the conjugate of the numerator \[\lim_{h\to 0}\frac{\sqrt{x+h+9}-\sqrt{x+9}}{h}\times\frac{\sqrt{x+h+9}+\sqrt{x+9}}{\sqrt{x+h+9}+\sqrt{x+9}}\]

OpenStudy (anonymous):

leave the denomiator in factored form. the numerator will just be \(h\) cancel the \(h\) top and bottom, then replace \(h\) by 0

OpenStudy (anonymous):

so after the conjugate multiplying in, it would be x+h+9 - x -9 / h(the sqrt plus the other sqrt) right?

OpenStudy (anonymous):

so 1/ the sqrt + the sqrt

OpenStudy (anonymous):

got it thanks